Calculate the change in heat when 15.00 g of water vapor (steam) at 100.0°C condenses to liquid water and then cools to 27.00 °C.
1
Expert's answer
2020-11-04T14:23:06-0500

E(condensing) = L * m = 2260 J/g * 15 g = 33900 J

E(cooling) = c*m*"\Delta"T = 4.2 J/(g*K) * 15 g * 73 K= 4599 J

change in heat = 33900 J + 4599 J = 38499 J = 38.5 kJ